OpenStudy (anonymous):

3x+y=12 y=-x-2 solve please?

jimthompson5910 (jim_thompson5910):

3x+y=12 3x+(-x-2)=12 3x-x-2=12 2x-2=12 2x=12+2 2x=14 x=14/2 x=7 -------------------------------- y=-x-2 y=-7-2 y=-9 So the solutions are x = 7 and y = -9